# Gap solitons of the Wannier and Bloch types in spin-orbit-coupled Bose-Einstein condensates with a moiré lattice

### v1

@@ III.2 Wannier-type and Bloch-type gap solitons
Figure 3: WT (Wannier-type) GS families bifurcating from the lowest two flat bands for the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$. Other parameters are $\gamma=0.5$, $c_{0}=1$ and $c_{2}=-1.5$ (the ferromagnetic system). (a)-(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)-(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. The drawing grid is $(x,y)\in[-3,3]$. (f) Norm NN versus the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The black lines represent two flat bands with $\mu_{1}=1.6998$ and $\mu_{2}=1.8065$.
Comparing the norm curves of these two GS families ($\varphi_{j}^{(1)}$ and $\varphi_{j}^{(2)}$) for different interaction types, antiferromagnetic and ferromagnetic (the red and green lines in Figs. 2(f) and 3(f)), we notice the difference that the GSs exist in the interval $\mu>\mu_{n}$ when $c_{0}>0$, i.e., $\Delta\mu>0$ due to the decisive role of $c_{0}$. Moreover, we also find three other GS types, which have different density profiles of components $\varphi_{+1}$ and $\varphi_{-1}$. In Fig. 3(c), the GS family with different center positions of three components reveals the phase separation, cf. Ref. [59]. The two degenerate GS families in Figs. 3(d) and 3(e) represent vortex states with a winding number difference of $1$ between the components [60]. These are “exotic” quantum states induced by SOC. For these GS families with $|\varphi_{+1}|^{2}\neq|\varphi_{-1}|^{2}$, the interaction terms with $c_{2}$ dominate over those with $c_{0}$, hence $c_{2}$ determines the offset of the chemical potential, producing $\Delta\mu<0$ ($c_{2}<0$), as indicated by the norm curves (the blue and magenta lines) in Fig. 3(f). In fact, the three GS families can be approximated by a linear combination of two Wannier functions $w_{j}^{(1)}$ and $w_{j}^{(2)}$, which are, essentially, the superpositions of two corresponding GS families (with the same $\mu$) in Figs. 3(a) and 3(b). Using numerical data, these relations can be written as
@@
Figure 4: BT (Bloch-type) GS families bifurcating from the lowest two energy bands when $\theta=\arctan(3/4)$, $a=0.5\pi$, $V_{0}=4$, $\gamma=0.5$, $c_{0}=1$, and $c_{2}=-1.5$. (a)-(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)-(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. Here, the drawing grid is $(x,y)\in[-4,4]$. (f) The norm NN versus the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The blue area represents the energy band, and the inset shows the magnified view of the region enclosed by the black dashed lines.
As the ML period and lattice depth decrease, the low-energy linear bands gradually widen, and the energy gaps between them get narrow, until they collapse, leaving solely the semi-infinite gap. For example, in the case of $a=0.5\pi$ and $V_{0}=4$, the lowest two energy bands both exhibit finite bandwidth (this, being non-flat). Based on the norm curves of several WT GS families for different interaction strengths, we can identify GS families bifurcating from the lowest two energy bands in the case of a small ML period ($a=0.5\pi$), whose phases are similar to those of WT GS families ($a=2\pi$). Five GS families for $c_{0}=1$ and $c_{2}=-1.5$ are shown in Figs. 4(a)-4(e). These GSs span multiple unit cells, with envelopes modulated by the ML. The corresponding norm curves exhibit discontinuity while they hit energy bands, accompanied by a significant broadening of the envelope when $\mu$ approaches the band edges, as shown in Fig. 4(f). This broadening effect intensifies with increasing bandwidth. The existence of a finite threshold, $N_{\mathrm{th}}=\min(N)\gg 0$, implies that these GSs cannot maintain localization and approach the Bloch states when the norm is very small. Therefore, these GSs do not admit the approximation by Wannier functions, while they bifurcate from the energy bands with finite bandwidth [58], and thus they are termed BT GSs. Nevertheless, these GS families still satisfy the relation similar to that given by Eq. (11). The relation helps to find other BT GS families.
@@ III.3 Stability analysis of the gap solitons
Stability is the crucial characteristic of 2D solitons. We have systematically investigated the stability of the GSs families in the framework of the linearized Bogoliubov-de Gennes equations for small perturbations [57]. To this end, small perturbations $a_{j}(x,y)$ and $b_{j}(x,y)$ were added to the GS solutions as
@@
where $L_{0}=\nabla^{2}/2+\mu-V$, $C_{+}=c_{0}+c_{2}$, and $C_{-}=c_{0}-c_{2}$. By means of the numerical solution, we have obtained eigenvalues $\lambda$ corresponding to each stationary GS solution. Taking into regard possible numerical errors, we define the GS to be stable provided that $\max\left(\mathrm{Re}(\lambda)\right)<10^{-3}$. The so established stability was verified through numerical simulations of the perturbed GSs. The stability of the above-found GS families is designated, respectively, by continuous and dashed segments of the corresponding norm curves in Figs. 2(f)-4(f). It is seen that the WT GSs are stable only when $N$ is sufficiently small. Note that, among the GS families that branch off to the left (right) from the bifurcation point, those with non-zero (zero) vorticity have larger stability domains. The BT GSs exhibit stability properties similar to those of their WT GSs counterparts. Note that the BT GSs are unstable as they approach edges of the energy bands, which is essential as the energy bands widen, in particular, with the reduction of the lattice depth $V_{0}$. We performed extensive numerical simulations by using numerical solutions with random perturbations added to the input. The results, which are consistent with the linear stability analysis, are exemplified by three cases corresponding to the GSs shown in Figs. 2(a)-4(a), as shown in Fig. 5.
Figure 5: Panels (a)-(c) show the evolution of the maximum densities for the three components as revealed by the numerical simulations of the GSs in Figs. 2(a)-4(a), respectively. The corresponding spatial density profiles of all the three components at t = 100 are displayed in panels (d)-(f).
@@ III.4 The transition between Wannier-type and Bloch-type gap solitons induced by spin-orbit coupling
Figure 6: (a) The effect of the SOC strength $\gamma$ on the flatness $\Delta_{1}$ of the lowest-energy band for $V_{0}=2$ (the red line) and $V_{0}=4$ (the black line). (b) The inverse participation ratio (IPR) of the GS solutions bifurcating from the lowest-energy band versus $\gamma$ for different values of NN when $V_{0}=2$. The corresponding subfigures show the density profiles of the component $\varphi_{+1}$ at $\gamma=1.3$ and $\gamma=2$. Panel (c) displays the corresponding norm curves and the lowest-energy bands for the respective values of $\gamma$ marked in (a). The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). Here, $\theta=\arctan(3/4)$, $a=0.5\pi$, $c_{0}=-1$, and $c_{2}=1.5$.
We find that increasing $\gamma$ leads to a periodic decrease and increase in the flatness of the lowest-energy band, but overall, $\Delta_{1}$ exhibits a decreasing trend.
To systematically investigate the effect of the SOC strength on GSs, we consider the GS solutions bifurcating from the lowest-energy band at a fixed norm for different values of $\gamma$. The results show that as $\gamma$ increases, the soliton profile narrows and its amplitude grows, indicating enhanced localization. This trend is clearly demonstrated by the inverse participation ratio (IPR),
@@
which quantifies the localization degree, and by direct comparison of the density distributions at $\gamma=1.3$ and $\gamma=2$, as shown in Fig. 6(b). The focusing effect induced by $\gamma$ is also the reason why SOC can stabilize two-dimensional solitons in the absence of an external potential. Furthermore, the norm curves of GS families are compared for $\gamma=0.5$, $\gamma=1.3$, $\gamma=1.45$, $\gamma=1.76$, and $\gamma=2$ in Fig. 6(c). These results show a distinct transition in the character of the norm curves: they evolve from non-monotonous to monotonous ones as $\gamma$ increases, which is accompanied by a reduction in the norm threshold. Both the change in monotonicity and the reduction of norm threshold consistently indicate a gradual transition of the soliton toward the Wannier type. Therefore, we conclude that, for appropriately selected lattice parameters, the increase in $\gamma$ can enhance the band flatness to the level necessary for the emergence of WT GSs, driving the BT $\rightarrow$ WT transition of GSs.
@@ III.5 Soliton families in quasiperiodic moiré lattices
For a Pythagorean angle $\theta^{\prime}$, the corresponding Pythagorean triple ($\bar{a},\bar{b},\bar{c}$) can be represented by two coprime natural numbers ($\bar{m},\bar{n}$), namely $\bar{a}=\bar{m}^{2}-\bar{n}^{2}$, $\bar{b}=2\bar{m}\bar{n}$, and $\bar{c}=\bar{m}^{2}+\bar{n}^{2}$. By solving $\sin\theta\approx\sin\theta^{\prime}=(\bar{m}^{2}-\bar{n}^{2})/(\bar{m}^{2}+\bar{n}^{2})$, all Pythagorean angles that approximate a given non-Pythagorean angle $\theta$ can be derived. For $\theta=\pi/4$, the two coprime natural numbers satisfy a ratio condition $\bar{m}/\bar{n}=\sqrt{3+2\sqrt{2}}\approx 2.41$. The approximate solutions ($\bar{m},\bar{n}$) include $(5,2)$, $(12,5)$, etc., and the corresponding Pythagorean triples ($\bar{a},\bar{b},\bar{c}$) are $(21,20,29)$ and ($119,120,169$), respectively. In Figs. 7(a)-7(c),
@@
the direct comparison of the ML profiles for the non-Pythagorean angle ($\theta=\pi/4$) and two Pythagorean angles ($\theta^{\prime}=\arctan(20/21)$ and $\theta^{\prime}=\arctan(119/120)$) clearly demonstrates the accuracy of the periodic approximation method.
Quasiperiodic potentials enable the formation of localized states in purely linear systems through the Anderson-localization mechanism [61]. A natural question is what happens to 2D solitons under the action of quasiperiodic potentials. Here, we address this issue in the case of $\theta=\pi/4$, not addressing the edge states induced by finite boundaries [62]. For the large lattice period $a=2\pi$, the quasiperiodic ML still supports all bright GS families shown in Figs. 2 and 3. Profiles of these solitons remain nearly identical to those in the periodic ML with $\theta=\arctan(3/4)$, and the corresponding norm curves nearly coincide. The point is that these solitons are essentially localized in the central parabolic-like potential when $a=2\pi$, rendering the effective models, and solitons produced by the models, nearly indistinguishable for Pythagorean and non-Pythagorean twist angles.
@@
Figure 8: GS families for the non-Pythagorean angle $\theta=\pi/4$ when $a=0.5\pi$, $V_{0}=4$, $\gamma=0.5$, $c_{0}=1$, and $c_{2}=-1.5$. (a)-(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)-(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. The drawing grid is $(x,y)\in[-4,4]$. (f) Norm NN versus the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The red dots represent the linear spectrum for the ML with $\theta=\pi/4$. The blue areas represent the energy bands for the ML with $\theta^{\prime}=\arctan(119/120)$.
Notably, the GSs supported by the ML with $\theta=\pi/4$ exhibit significantly different density distributions from those at $\theta=\arctan(3/4)$, but they still satisfy the approximate linear relation similar to that described in Eq. (11). The numerically generates norm curves of five GS families are plotted in Fig. 8(f), including their stability. The norm curves for soliton families in Figs. 8(c)-8(e) maintain continuity, while the curves for soliton families in Figs. 8(a) and 8(b) exhibit two discontinuities (where the soliton does not exist), indicating the presence of continuous energy spectra (without gaps) in the two regions. To verify this, we directly compute the linear spectrum (red dots) of the quasiperiodic system at $\theta=\pi/4$ alongside the bandgap structure (blue areas) of the periodic system at $\theta^{\prime}=\arctan(119/120)$, as shown in Fig. 8(f). In the linear spectrum, the two isolated red dots on the left precisely locate bifurcation points of the soliton families, as do the two flat bands in the bandgap structure. Moreover, regions with dense sets of eigenvalues in the linear spectrum and broad bands in the bandgap structure clearly display the discontinuous parts of the norm curves, where solitons are absent. These inferences further validate the accuracy of the norm curves obtained by the numerical method. The comparison of Figs. 4(f) and 8(f) reveals that the stability is primarily governed by the type of the soliton, independent of twist angles. On the other hand, we find SOC can also enhance the localization of the solitons in quasiperiodic MLs.

### v2

@@ III.2 Wannier-type and Bloch-type gap solitons
Figure 3: WT (Wannier-type) GS families bifurcating from the lowest two flat bands for the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$. Other parameters are $\gamma=0.5$, $c_{0}=1$, and $c_{2}=-1.5$ (the ferromagnetic system). (a)–(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)–(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. The drawing grid is $(x,y)\in[-3,3]$. (f) Norm NN vs the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The black lines represent two flat bands with $\mu_{1}=1.6998$ and $\mu_{2}=1.8065$.
Comparing the norm curves of these two GS families $\varphi_{j}^{(1)}$ and $\varphi_{j}^{(2)}$] for different interaction types, antiferromagnetic and ferromagnetic [the red and green lines in Figs. 2(f) and 3(f)], we notice the difference that the GSs exist in the interval $\mu>\mu_{n}$ when $c_{0}>0$, i.e., $\Delta\mu>0$ due to the decisive role of $c_{0}$. Moreover, we also find three other GS types, which have different density profiles of components $\varphi_{+1}$ and $\varphi_{-1}$. In Fig. 3(c) the GS family with different center positions of three components reveals the phase separation, cf. Ref. [59]. The two degenerate GS families in Figs. 3(d) and 3(e) represent vortex states with a winding number difference of $1$ between the components [60]. These are “exotic” quantum states induced by SOC. For these GS families with $|\varphi_{+1}|^{2}\neq|\varphi_{-1}|^{2}$, the interaction terms with $c_{2}$ dominate over those with $c_{0}$; hence $c_{2}$ determines the offset of the chemical potential, producing $\Delta\mu<0$ ($c_{2}<0$), as indicated by the norm curves (the blue and magenta lines) in Fig. 3(f). In fact, the three GS families can be approximated by a linear combination of two Wannier functions $w_{j}^{(1)}$ and $w_{j}^{(2)}$, which are, essentially, the superpositions of two corresponding GS families (with the same $\mu$) in Figs. 3(a) and 3(b). Using numerical data, these relations can be written as
@@
Figure 4: BT (Bloch-type) GS families bifurcating from the lowest two energy bands when $\theta=\arctan(3/4)$, $a=0.5\pi$, $V_{0}=4$, $\gamma=0.5$, $c_{0}=1$, and $c_{2}=-1.5$. (a)–(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)–(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. Here, the drawing grid is $(x,y)\in[-4,4]$. (f) The norm NN vs the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The blue area represents the energy band, and the inset shows the magnified view of the region enclosed by the black dashed lines.
As the ML period and lattice depth decrease, the low-energy linear bands gradually widen, and the energy gaps between them get narrow, until they collapse, leaving solely the semi-infinite gap. For example, in the case of $a=0.5\pi$ and $V_{0}=4$, the lowest two energy bands both exhibit finite bandwidth (this, being non-flat). Based on the norm curves of several WT GS families for different interaction strengths, we can identify GS families bifurcating from the lowest two energy bands in the case of a small ML period ($a=0.5\pi$), whose phases are similar to those of WT GS families ($a=2\pi$). Five GS families for $c_{0}=1$ and $c_{2}=-1.5$ are shown in Figs. 4(a)-4(e). These GSs span multiple unit cells, with envelopes modulated by the ML. The corresponding norm curves exhibit discontinuity while they hit energy bands, accompanied by a significant broadening of the envelope when $\mu$ approaches the band edges, as shown in Fig. 4(f). This broadening effect intensifies with increasing bandwidth. The existence of a finite threshold, $N_{\mathrm{th}}=\min(N)\gg 0$, implies that these GSs cannot maintain localization and approach the Bloch states when the norm is very small. Therefore, these GSs do not admit the approximation by Wannier functions, while they bifurcate from the energy bands with finite bandwidth [58] and thus they are termed BT GSs. Nevertheless, these GS families still satisfy the relation similar to that given by Eq. (11). The relation helps to find other BT GS families.
@@ III.3 Stability analysis of the gap solitons
Stability is the crucial characteristic of 2D solitons. We have systematically investigated the stability of the GSs families in the framework of the linearized Bogoliubov–de Gennes equations for small perturbations [57]. To this end, small perturbations $a_{j}(x,y)$ and $b_{j}(x,y)$ were added to the GS solutions as
@@
where $L_{0}=\nabla^{2}/2+\mu-V$, $C_{+}=c_{0}+c_{2}$, and $C_{-}=c_{0}-c_{2}$. By means of the numerical solution, we have obtained eigenvalues $\lambda$ corresponding to each stationary GS solution. Taking into account possible numerical errors, we define the GS to be stable provided that $\max\left(\mathrm{Re}(\lambda)\right)<10^{-3}$. The so established stability was verified through numerical simulations of the perturbed GSs. The stability of the above-found GS families is designated, respectively, by continuous and dashed segments of the corresponding norm curves in Figs. 2(f)-4(f). It is seen that the WT GSs are stable only when $N$ is sufficiently small. Note that among the GS families that branch off to the left (right) from the bifurcation point, those with non zero (zero) vorticity have larger stability domains. The BT GSs exhibit stability properties similar to those of their WT GSs counterparts. Note that the BT GSs are unstable as they approach edges of the energy bands, which is essential as the energy bands widen, in particular, with the reduction of the lattice depth $V_{0}$. We performed extensive numerical simulations by using numerical solutions with random perturbations added to the input. The results, which are consistent with the linear stability analysis, are exemplified by three cases corresponding to the GSs shown in Figs. 2(a)-4(a), as shown in Fig. 5.
Figure 5: Panels (a)–(c) show the evolution of the maximum densities for the three components as revealed by the numerical simulations of the GSs in Figs. 2(a)-4(a), respectively. The corresponding spatial density profiles of all three components at $t=100$ are displayed in panels (d)–(f).
@@ III.4 The transition between Wannier-type and Bloch-type gap solitons induced by spin-orbit coupling
Figure 6: (a) The effect of the SOC strength $\gamma$ on the flatness $\Delta_{1}$ of the lowest-energy band for $V_{0}=2$ (the red line) and $V_{0}=4$ (the black line). (b) The inverse participation ratio (IPR) of the GS solutions bifurcating from the lowest-energy band vs $\gamma$ for different values of NN when $V_{0}=2$. The corresponding subfigures show the density profiles of the component $\varphi_{+1}$ at $\gamma=1.3$ and $\gamma=2$. Panel (c) displays the corresponding norm curves and the lowest-energy bands for the respective values of $\gamma$ marked in (a). The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). Here, $\theta=\arctan(3/4)$, $a=0.5\pi$, $c_{0}=-1$, and $c_{2}=1.5$.
We find that increasing $\gamma$ leads to a periodic decrease and increase in the flatness of the lowest-energy band, but overall, $\Delta_{1}$ exhibits a decreasing trend. To systematically investigate the effect of the SOC strength on GSs, we obtain the GS solutions bifurcating from the lowest-energy band at a fixed norm for different values of $\gamma$. The results show that as $\gamma$ increases, the soliton profile narrows and its amplitude grows, indicating enhanced localization. This trend is clearly reflected both in the inverse participation ratio (IPR),
@@
which quantifies the degree of localization, and in a direct comparison of the density distributions at $\gamma=1.3$ and $\gamma=2$, as shown in Fig. 6(b). The focusing effect induced by $\gamma$ is also the reason why SOC can stabilize two-dimensional solitons in the absence of an external potential. Furthermore, the norm curves of GS families are compared for $\gamma=0.5$, $\gamma=1.3$, $\gamma=1.45$, $\gamma=1.76$, and $\gamma=2$ in Fig. 6(c). These results show a distinct transition in the character of the norm curves: the monotonicity evolves from non monotonic to monotonic as $\gamma$ increases, accompanied by a reduction in the norm threshold. Both the change in monotonicity and the reduction in norm threshold consistently indicate a gradual transition of the soliton type toward the Wannier type. Therefore, we conclude that for appropriately selected lattice parameters, an increase in $\gamma$ can enhance the band flatness to the level necessary for the emergence of WT GSs, driving the transition of BT GSs to WT GSs.
@@ III.5 Soliton families in quasiperiodic moiré lattices
For a Pythagorean angle $\theta^{\prime}$, the corresponding Pythagorean triple ($\bar{a},\bar{b},\bar{c}$) can be represented by two coprime natural numbers ($\bar{m},\bar{n}$), namely, $\bar{a}=\bar{m}^{2}-\bar{n}^{2}$, $\bar{b}=2\bar{m}\bar{n}$, and $\bar{c}=\bar{m}^{2}+\bar{n}^{2}$. By solving $\sin\theta\approx\sin\theta^{\prime}=(\bar{m}^{2}-\bar{n}^{2})/(\bar{m}^{2}+\bar{n}^{2})$, all Pythagorean angles that approximate a given non-Pythagorean angle $\theta$ can be derived. For $\theta=\pi/4$, the two coprime natural numbers satisfy a ratio condition $\bar{m}/\bar{n}=\sqrt{3+2\sqrt{2}}\approx 2.41$. The approximate solutions ($\bar{m},\bar{n}$) include $(5,2)$, $(12,5)$, etc., and the corresponding Pythagorean triples ($\bar{a},\bar{b},\bar{c}$) are $(21,20,29)$ and ($119,120,169$), respectively. In Figs. 7(a)-7(c),
@@
the direct comparison of the ML profiles for the non-Pythagorean angle ($\theta=\pi/4$) and two Pythagorean angles $\theta^{\prime}=\arctan(20/21)$ and $\theta^{\prime}=\arctan(119/120)$] clearly demonstrates the accuracy of the periodic approximation method.
Quasiperiodic potentials enable the formation of localized states in purely linear systems through the Anderson-localization mechanism [61]. A natural question is what happens to 2D solitons under the action of quasiperiodic potentials. Here, we address this issue in the case of $\theta=\pi/4$, not considering the edge states induced by finite boundaries [62]. For the large lattice period $a=2\pi$, the quasiperiodic ML still supports all bright GS families shown in Figs. 2 and 3. Profiles of these solitons remain nearly identical to those in the periodic ML with $\theta=\arctan(3/4)$, and the corresponding norm curves nearly coincide. The point is that these solitons are essentially localized in the central parabolic-like potential when $a=2\pi$, rendering the effective models, and solitons produced by the models, nearly indistinguishable for Pythagorean and non-Pythagorean twist angles.
@@
Figure 8: GS families for the non-Pythagorean angle $\theta=\pi/4$ when $a=0.5\pi$, $V_{0}=4$, $\gamma=0.5$, $c_{0}=1$, and $c_{2}=-1.5$. (a)–(e) The density profiles of three components $\varphi_{+1}$ (the first row), $\varphi_{-1}$ (the second row), and $\varphi_{0}$ (the third row) for five GS families. The corresponding phases are shown in the subfigures. The first and second rows in (a)–(c) share the same color bars. $|\varphi_{-1}|$ in (d) and (e) are equal to $|\varphi_{+1}|$ in (e) and (d), respectively. The drawing grid is $(x,y)\in[-4,4]$. (f) Norm NN vs the chemical potential $\mu$ for the five GS families. The norm curves for (d) and (e) are identical. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The red dots represent the linear spectrum for the ML with $\theta=\pi/4$. The blue areas represent the energy bands for the ML with $\theta^{\prime}=\arctan(119/120)$.
Notably, the GSs supported by the ML with $\theta=\pi/4$ exhibit significantly different density distributions from those at $\theta=\arctan(3/4)$, but they still satisfy the approximate linear relation similar to that described in Eq. (11). The numerically generates norm curves of five GS families are plotted in Fig. 8(f), including their stability. The norm curves for soliton families in Figs. 8(c)-8(e) maintain continuity, while the curves for soliton families in Figs. 8(a) and 8(b) exhibit two discontinuities (where the soliton does not exist), indicating the presence of continuous energy spectra (without gaps) in the two regions. To verify this, we directly compute the linear spectrum (red dots) of the quasiperiodic system at $\theta=\pi/4$ alongside the band-gap structure (blue areas) of the periodic system at $\theta^{\prime}=\arctan(119/120)$, as shown in Fig. 8(f). In the linear spectrum, the two isolated red dots on the left precisely locate bifurcation points of the soliton families, as do the two flat bands in the band-gap structure. Moreover, regions with dense sets of eigenvalues in the linear spectrum and broad bands in the band-gap structure clearly display the discontinuous parts of the norm curves, where solitons are absent. These inferences further validate the accuracy of the norm curves obtained by the numerical method. The comparison of Figs. 4(f) and 8(f) reveals that the stability is primarily governed by the type of the soliton, independent of twist angles. On the other hand, we find SOC can also enhance the localization of the solitons in quasiperiodic MLs.